OpenStudy (anonymous):

f(x)=√(4-x²), 0 ≤ x ≤ 2 Find the inverse

OpenStudy (valpey):

\[f(x) = \sqrt{4-x^2} = y\] \[4-x^2 = y^2\] \[-x^2 =-4+ y^2\] \[x^2 =4- y^2\] \[ \sqrt{x^2} = \sqrt{4- y^2}\] \[ x = \sqrt{4- y^2} = g(y)\] g is the inverse of f on \[0\le y\le2\]